Factorise the following expressions
 a^4 +2a^2 b^2 +b^4

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

a^4+2a^2b2+b^4
= (a^2)^2+2×a^2×b^2+(b^2)^2
= (a^2+b^2)^2
Using identity: (x+y)^2= x^2+2xy+y^2

Thanks